OpenStudy (anonymous):

what is 1divided by x minus 1 divided by y divided by 1 divided by x

OpenStudy (anonymous):

well i guess, 1/x-x/y. ryt?

OpenStudy (anonymous):

According to WolframAlpha, "1divided by x minus 1 divided by y divided by 1 divided by x" is:\[\frac{1}{x}-\frac{1}{x y} \]

OpenStudy (anonymous):

it is if we follow the same predence operators from left to right.. for example: 1/x-(1/y)/1/x the ur ans comes.. bu if we take it like this: 1/x-(1/y)/(1/x) the my ans comes.. strange.. so are we supposed to go left to right with same precedence?

OpenStudy (anonymous):

ok what if i had a problem that was 3/x^2-5x+6-2/x^2-9

OpenStudy (anonymous):

this is so confusing

OpenStudy (anonymous):

just go with the BODMAS rule.

OpenStudy (anonymous):

One major problem with the problem statements on this site is this: What does a+b/c mean? Is it a + (b/c) or (a+b)/c ? In the programming languages of the computer world it is the former because by general agreement, the multiplication and division operators, * and / respectively, have higher precedence, binding power so to speak, than addition and subtraction, + and - . Rewriting the problem expression using the precedence rules above, I got: (1divided by x )minus( ((1divided by y )divided by 1) divided by x) When the modified expression above was fed to WolframAlpha, they responded with:\[\frac{1}{x}-\frac{1}{x y} \]the same previous answer. The attachement, test.jpg, is a partial screen capture of a Mathematica 8 Notebook showing the second submission to WolframAlpha. Mathematica 8 notebooks have a direct connection with WolframAlpha.